Parker Rogers will turn 30 years old tomorrow. He comes up with a plan to save for his retirement at 65 years of age. He has set himself a retirement target of $2,000,000. He currently has an inheritance of $50,000 invested in a money market account earning 4.5 percent. He plans to leave the money market account as part of his retirement savings portfolio. To achieve his retirement goal, he plans to put aside an extra savings every year, starting next year (at his 31st birthday), in a mutual fund that will earn 9.5 percent annually. How much will Gary have to save every year?

How would you work this problem using a financial calculator ?
Answer #1

Amount required = FV = 2000,000

Future value of $50,000 in money market account = PV*(1+r)^n

= 50000*1.045^35

=233367.40

Balance requirement = 2000000-233367.40 = 1766632.60

Using financial calculator
Input:

FV = 1766632.60

I/Y = 9.5

N = 65-30 = 35

Solve for PMT = -7309.54

Gary will have to save $7309.54 every year